OpenStudy (anonymous):

How do i tell if this trinomial 8x^2 - 14x-3 is prime or not?

OpenStudy (anonymous):

Is it prime over the integers, the rationals or the real numbers?

OpenStudy (anonymous):

your equation \[ 8 x^2-14 x-3=0 \] Has two real roost but they are not rationals. Here they are \[ \left\{\left\{x= \frac{1}{8} \left(7-\sqrt{73}\right)\right\},\left\{x= \frac{1}{8} \left(7+\sqrt{73}\right)\right\}\right\} \] So your polynomial is prime over the rationals, but not prime over the reals.